OpenStudy (anonymous):

help please Evaluate S5 for 400 + 200 + 100 + … and select the correct answer below. 25 775 1,125 500

OpenStudy (welshfella):

you can just find what the 4 th and 5th terms are then add them up Do you see the pattern?

OpenStudy (welshfella):

how do you get from 400 to 200 and from 200 to 100? what process is done on 400 and 200 to get the next number?

OpenStudy (welshfella):

and what would you do to 100 to get the next number in the series?

OpenStudy (anonymous):

would the answer be 25?

OpenStudy (welshfella):

No how could it be 25? S5 means the Sum of 5 terms

OpenStudy (anonymous):

Oh i don't know. so what would the answer be

OpenStudy (welshfella):

to get the next term in the series you multiply by 0.5 400 * 0.5 = 200 200 * 0.5 = 100 so what are the next 2 terms?

OpenStudy (welshfella):

0.5 * 100 = ?

OpenStudy (anonymous):

100 * 0.5 = 50 50 * 0.5 = 25

OpenStudy (welshfella):

exactly so you have the first 5 terms so all you have to do is add them up 400+ 200 etc

OpenStudy (welshfella):

can you do that S5 = 400+200+100+50+25

OpenStudy (anonymous):

yes i got it thank you

OpenStudy (welshfella):

yw